# Unpacking Let Alone: Human-Scale Models Generalize to a Rare Construction in Form but not Meaning

### v1

@@ 8 Conclusion
This work has shown that BabyLM models can learn various formal properties of let-alone. This formal competence is maintained even in the absence of direct evidence of let-alone and related constructions. Such a result points to the crucial nature of indirect evidence for learning a construction as rare as let-alone. We additionally find that the individual tokens of let and alone, independent of their occurrence in the construction, are crucial to performance overall, raising the question of how closely linked these lexical items are to let-alone as a construction.
On the other hand, BabyLM models seem to have no grasp of let-alone’s meaning. This result underscores the importance of considering both formal and functional competence when assessing LM capabilities regarding a specific linguistic phenomenon, and doing so in a controlled manner. While the formal capabilities of BabyLM models are promising, insofar as we consider meaning to be a central part of language, our results cast doubt on the proposition that robust semantics—including of rare constructions—can be learned from form alone via human-scale pretraining.
9 Limitations
@@ Appendix A Pretraining Hyperparameters
Table 5 reports the hyperparameters for BabyLM pretraining, which were kept consistent across all runs. This is the same as those reported in Misra and Mahowald (2024), except that they tune the learning rate between the options of 1e-4, 3e-4, 1e-3, and 3e-3. We opt for 1e-4 after preliminary runs showed that higher learning rates led to development loss bottoming out much before the end of 20 epochs of training. In total, 12 complete pretraining runs were performed, totaling roughly 100 hours runtime on a NVIDIA A100 GPU.
@@ Appendix B GPT Prompting Experiment
For the semantic experiment (see section 5), we reformulated our task as a prompt-based experiment for input into GPT-4.1. Since this result serves primarily as a comparison point to BabyLM models, we do not perform prompt optimization. Thus, the reported GPT-4.1 accuracy should not be taken as an upper limit on LLM performance on this dataset. GPT-4.1 was accessed through the OpenAI API. The cost of this experiment was roughly $8 USD. Table 6 reports the prompt used for GPT-4.1 experiments.
| Prompt template fed to GPT-4.1 |
@@
Appendix C Potential Risks
We do not foresee any potential risks that are associated with our work.

### v2

@@ 8 Conclusion
This work has shown that BabyLM models can learn various formal properties of let-alone. This formal competence is maintained even in the absence of direct evidence of let-alone and related constructions. Such a result points to the crucial nature of indirect evidence for learning a construction as rare as let-alone. On the other hand, BabyLM models seem to have very little grasp of let-alone’s meaning. This result underscores the importance of considering both formal and functional competence when assessing LM capabilities regarding a specific linguistic phenomenon, and doing so in a controlled manner. While the formal capabilities of BabyLM models are promising, insofar as we consider meaning to be a central part of language, our results cast doubt on the proposition that robust semantics—including of rare constructions—can be learned from form alone via human-scale pretraining.
Limitations
@@
Ethics and Risks
We do not foresee any major ethical concerns or risks associated with this work. All evaluation data is templatically created and thus free of any sensitive or offensive content. The dataset released here is designed for research into the let-alone construction, and the authors do not foresee that it could be deployed for any malicious purpose.
Acknowledgments
We thank the anonymous reviewers, ACs, and SACs for EMNLP and the ARR May Cycle. We also thank members of the NERT and PiCoL labs for their useful comments on early versions of this work. This research was supported in part by NSF award IIS-2144881.
@@ Appendix A Pretraining Hyperparameters
Table˜5 reports the hyperparameters for BabyLM pretraining, which were kept consistent across all runs. This is the same as those reported in Misra and Mahowald (2024), except that they tune the learning rate between the options of 1e-4, 3e-4, 1e-3, and 3e-3. We opt for 1e-4 after preliminary runs showed that higher learning rates led to development loss bottoming out much before the end of 20 epochs of training. In total, 12 complete pretraining runs were performed, totaling roughly 100 hours runtime on a NVIDIA A100 GPU.
@@ Appendix B GPT Prompting Experiment
For the semantic experiment (see section˜5), we reformulated our task as a prompt-based experiment for input into GPT-4.1. Since this result serves primarily as a comparison point to BabyLM models, we do not perform prompt optimization. Thus, the reported GPT-4.1 accuracy should not be taken as an upper limit on LLM performance on this dataset. GPT-4.1 was accessed through the OpenAI API. The cost of this experiment was roughly $8 USD. Table˜6 reports the prompt used for GPT-4.1 experiments.
| Prompt template fed to GPT‑4.1 |